# Modular Machine Learning: An Indispensable Path towards New-Generation Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have dramatically advanced machine learning research including natural language processing, computer vision, data mining, etc., yet they still exhibit critical limitations in reasoning, factual consistency, and interpretability. In this paper, we introduce a novel learning paradigm—Modular Machine Learning (MML)—as an essential approach toward new-generation LLMs. MML decomposes the complex structure of LLMs into three interdependent components: modular representation, modular model, and modular reasoning, aiming to enhance LLMs’ capability of counterfactual reasoning, mitigating hallucinations, as well as promoting fairness, safety, and transparency. Specifically, the proposed MML paradigm can: i) clarify the internal working mechanism of LLMs through the disentanglement of semantic components; ii) allow for flexible and task-adaptive model design; iii) enable interpretable and logic-driven decision-making process. We present a feasible implementation of MML-based LLMs via leveraging advanced techniques such as disentangled representation learning, neural architecture search and neuro-symbolic learning. We critically identify key challenges, such as the integration of continuous neural and discrete symbolic processes, joint optimization, and computational scalability, present promising future research directions that deserve further exploration. Ultimately, the integration of the MML paradigm with LLMs has the potential to bridge the gap between statistical (deep) learning and formal (logical) reasoning, thereby paving the way for robust, adaptable, and trustworthy AI systems across a wide range of real-world applications.
@@
The advent of Large Language Models (LLMs) [1], epitomized by the likes of ChatGPT, has undeniably been a watershed moment in the evolution of the AI landscape. These models have astounded the research community and the general public alike with their seemingly superhuman language-processing capabilities. In a plethora of tasks, they have managed to mimic human-level language understanding and generation with remarkable fidelity. From drafting eloquent essays and engaging in fluent conversations to providing detailed summaries of complex texts, LLMs have proven their capabilities across various scenarios. This remarkable performance has led some researchers to advocate the “bigger is better” principle. The underlying rationale is that as the model size and volume of training data grow exponentially (i.e., the scaling law), so does the breadth and depth of knowledge that the LLM can encapsulate. This enables the model to handle a wide range of language-related challenges with an accuracy and fluency that were previously unimaginable.
@@
However, as with any technological marvel, LLMs have their own Achilles’ heels. A major limitation of LLMs surfaces when it comes to quantitative reasoning tasks. For instance, LLMs often struggle with simple two-digit arithmetic problems. Without human interaction, they have trouble applying basic mathematical principles to obtain the right answers. This deficiency highlights the fact that, despite their remarkable linguistic capabilities, their cognitive and logical reasoning abilities, particularly in domains requiring formal and rule-based understanding, remain in need of substantial enhancement. Consequently, the research community has witnessed a growing trend for augmenting LLMs with additional tools, where the additional tools are designed to fill the gaps in reasoning capabilities. For example, coupling an LLM with a symbolic solver can enhance LLM’s ability to handle mathematical and logical problems. By integrating external knowledge bases, LLMs can access domain-specific information that might not have been part of their original training data as well, thereby enhancing their adaptable problem-solving capabilities. Such hybrid systems could potentially release the full potential of LLMs, allowing them to overcome current limitations and become more versatile and reliable in a wide range of AI applications.
@@
LLMs are expected to be explainable, reliable, adaptable, and extendable in real-world applications. For example, in many critical domains such as healthcare, finance, and legal systems, the decisions made by LLMs can have far-reaching consequences. When a model provides a diagnosis or financial advice, patients and stakeholders need to understand the underlying rationale. Without explainability, it becomes difficult to validate the accuracy and reliability of the output. In a medical context, if an LLM suggests a particular treatment plan, both doctors and patients must clearly understand the reasoning behind that recommendation. This not only fosters trust but also enables error detection and correction. Furthermore, as the demands of applications evolve and new tasks emerge, LLMs need to be adaptable to unseen scenarios via progressively incorporating new knowledge, domains, and functions. In the fast-paced world of technology, a static LLM would quickly become obsolete and outdated. Additionally, as new scientific discoveries arise, an extensible LLM should be capable of incorporating these advancements to enhance its capabilities.
In this paper, we propose a novel learning paradigm, i.e., Modular Machine Learning (MML), which includes i) modular representation, ii) modular model and iii) modular reasoning for deep learning architectures such as LLMs, presenting a viable path to attaining the above crucial characteristics required in a wide range of real-world applications. For ease of understanding, we utilize a visual question answering (VQA) task as an example to illustrate the process of MML for LLMs in Fig. 1. To further validate the methodology, in Fig. 2 we provide an instantiation of MML under VQA task with a feasible implementation. 1) Disentangled Representation Learning (DRL) for modular representation allows for a more organized and interpretable structure of information within the LLMs. By disentangling complex representations into independent semantic dimensions, it becomes easier to understand and control the model’s internal working process, thus enhancing explainability. 2) Neural Architecture Search (NAS) for modular model enables the creation of modularized neural architectures with various modules in charge of different functions. Different modules can be optimized with respect to specific tasks, making the overall LLM capable of reliably completing complex tasks with an adaptive combination of different prerequisite tasks. The modularized architecture also facilitates parameter maintenance and updating within different modules, contributing to easy extension with new modules regarding new functions. 3) Neuro-Symbolic Learning (NSL) for modular reasoning bridges the gap between the black-box neural network inference process and human-understandable symbolic reasoning. The modularized design formalizes the inference steps, improving explainability and reliability as the decision-making process becomes more transparent and auditable, as well as promoting fairness via enabling the detection and correction of potential biases in the reasoning chain. As such, modular machine learning holds the potential to transform LLMs into more powerful, trustworthy, and versatile tools for real-world applications. More discussions regarding the implementation will be presented in Section 4.
We will first present the methodology of MML in Section 2, then discuss the significance of MML for LLMs in Section 3, followed by a detailed description of a feasible implementation of MML in Section 4. Additionally, we will also explore the future directions in Section 5. In sum, we aim to provide a comprehensive understanding of the potential and limitations of current machine learning techniques, with the hope that insights from this paper may contribute to the advancement of Artificial General Intelligence (AGI).

### v2

@@ Abstract
Large language models (LLMs) have substantially advanced machine learning research, including natural language processing, computer vision, data mining, etc., yet they still exhibit critical limitations in explainability, reliability, adaptability, and extensibility. In this paper, we overview a promising learning paradigm, i.e., Modular Machine Learning (MML), as an essential approach toward new-generation LLMs capable of addressing these issues. We begin by systematically and comprehensively surveying the existing literature on modular machine learning, with a particular focus on modular data representation and modular models. Then, we propose a unified MML framework for LLMs, which decomposes the complex structure of LLMs into three interdependent components: modular representation, modular model, and modular reasoning. Specifically, the MML paradigm discussed in this article is able to: i) clarify the internal working mechanism of LLMs through the disentanglement of semantic components; ii) allow for flexible and task-adaptive model design; iii) enable an interpretable and logic-driven decision-making process. We further elaborate a feasible implementation of MML-based LLMs via leveraging advanced techniques such as disentangled representation learning, neural architecture search and neuro-symbolic learning. Last but not least, we critically identify the remaining key challenges, such as the integration of continuous neural and discrete symbolic processes, joint optimization, and computational scalability, present promising future research directions that deserve further exploration. Ultimately, we believe the integration of the MML with LLMs has the potential to bridge the gap between statistical (deep) learning and formal (logical) reasoning, thereby paving the way for robust, adaptable, and trustworthy AI systems across a wide range of real-world applications.
@@
The advent of Large Language Models (LLMs) [1], epitomized by the likes of ChatGPT [2], has undeniably been a watershed moment in the evolution of the AI landscape. These models have astounded the research community and the general public alike with their seemingly superhuman language-processing capabilities. In a plethora of tasks, they have managed to mimic human-level language understanding and generation with remarkable fidelity. From drafting eloquent essays and engaging in fluent conversations to providing detailed summaries of complex texts, LLMs have proven their capabilities across various scenarios. This remarkable performance has led some researchers to advocate the “bigger is better” principle. The underlying rationale is that as the model size and volume of training data grow exponentially (i.e., the scaling law), so does the breadth and depth of knowledge that the LLM can encapsulate. This enables the model to handle a wide range of language-related challenges with an accuracy and fluency that were previously unimaginable.
@@
However, as with any technological marvel, LLMs have their own Achilles’ heels. A major limitation of LLMs surfaces when it comes to quantitative reasoning tasks. For instance, LLMs often struggle with simple two-digit arithmetic problems. Without human interaction, they have trouble applying basic mathematical principles to obtain the right answers. This deficiency highlights the fact that, despite their remarkable linguistic capabilities, their cognitive and logical reasoning abilities, particularly in domains requiring formal and rule-based understanding, remain in need of substantial enhancement. Consequently, the research community has witnessed a growing trend for augmenting LLMs with additional tools, where the additional tools are designed to fill the gaps in reasoning capabilities. For example, coupling an LLM with a symbolic solver can enhance the LLM’s ability to handle mathematical and logical problems. By integrating external knowledge bases, LLMs can access domain-specific information that might not have been part of their original training data as well, thereby enhancing their adaptable problem-solving capabilities. Such hybrid systems could potentially release the full potential of LLMs, allowing them to overcome current limitations and become more versatile and reliable in a wide range of AI applications.
@@
LLMs are expected to be explainable, reliable, adaptable, and extendable in real-world applications. In many critical domains such as healthcare, finance, and legal systems, the decisions made by LLMs can have far-reaching consequences. For example, when a model provides a diagnosis or financial advice, patients and stakeholders need to understand the underlying rationale. Without explainability, it becomes difficult to validate the accuracy and reliability of the output. Similarly, adaptability is essential as the demands of applications evolve and new tasks emerge; a static LLM would quickly become obsolete. Finally, extensibility is required to keep pace with new scientific discoveries and technological advancements, ensuring that models can incorporate and leverage new knowledge effectively.
Modular Machine Learning (MML) serves as one indispensable path to achieve the above properties for LLM. By decomposing complex systems into well-defined and interoperable modules, MML directly supports explainability through clearer structure and interpretability, being capable of tracing how specific outputs are produced [3, 4]. It also enhances reliability by allowing modules to be validated and improved independently, reducing the risk of systemic errors. MML fosters adaptability as well, since individual modules can be modified or replaced to meet new requirements without redesigning the entire system [5]. Finally, MML is able to promote extensibility, as new scientific methods or domain-specific knowledge can be incorporated as additional modules that integrate seamlessly into the existing architecture [6]. Unlike monolithic end-to-end approaches, the modular perspective provides a structured framework that enhances scalability, transparency, and control—qualities essential for deploying LLMs in real-world, high-stakes environments [7, 8, 9].
In this paper, we present a systematic and comprehensive survey of Modular Machine Learning (MML), which, to the best of our knowledge, is the first literature review of this paradigm. We begin with a formal definition of MML to clarify its conceptual position. We then categorize existing methods into two major groups: (i) modular data representation, which aims to separate complex observations into semantically meaningful and independent factors, and (ii) modular model optimization, which emphasizes decomposing architectures and reasoning processes into functional modules. Within each group, we review representative approaches under both supervised and unsupervised settings, highlighting their fundamental assumptions, technical designs, and trade-offs. Building upon this taxonomy, we further propose a unified MML framework for LLMs, consisting of three interdependent components: modular representation, modular model, and modular reasoning. For illustrative purposes, we employ the visual question answering (VQA) task as a running example (Fig. 1), which demonstrates how MML disentangles visual content, assembles specialized neural modules, and integrates symbolic reasoning to boost the ability of LLMs in generating interpretable outputs. To further contextualize the unified MML framework, we provide an instantiation of MML under the VQA task with a feasible implementation, where the three components are implemented as follows: (1) Disentangled Representation Learning (DRL) for modular representation enables a more organized and interpretable structure of information within LLMs by separating complex features into independent semantic dimensions, thereby enhancing transparency and controllability. (2) Neural Architecture Search (NAS) for modular model automates the construction of modularized neural architectures, where different modules are optimized for distinct functions and can be flexibly extended to new tasks. (3) Neuro-Symbolic Learning (NSL) for modular reasoning integrates symbolic rules with neural inference, thereby formalizing decision-making steps, improving reliability, and facilitating fairness by detecting and correcting potential biases. More detailed discussions of implementation aspects will be presented in Section 6. In general, this article provides a consolidated view of the field, i) overviewing existing literature comprehensively, and ii) positioning MML as a promising direction for enhancing the explainability, reliability, adaptability, and extensibility of large language models and related AI systems.
Relations to Relevant Surveys. Recently, several efforts have focused on topics that are related to a subtopic within a particular component of MML. Concretely, Disentangled Representation Learning (DRL) has been systematically reviewed as a paradigm for separating latent factors into independent and semantically meaningful components [10, 11], while Neural Architecture Search (NAS) has been extensively surveyed as an automated framework for discovering effective neural architectures under diverse constraints [12, 13, 14, 15, 16]. However, there have been no works that comprehensively discuss key elements such as Neuro-Symbolic Learning (NSL) and Modular Networks etc., not to mention the whole literature on MML. This paper goes beyond such domain-specific perspectives via providing a unified view that covers the complete collection of works in MML. Moreover, this work examines DRL and NAS from a Modular Perspective, which are totally different from those in relevant surveys. Last but not least, we propose a unified MML framework for LLMs with feasible implementations in this paper, which has huge potential in reshaping the design principles of next-generation large language models.
In summary, this survey aims to offer a consolidated understanding of the potential and limitations of existing approaches, with the goal of informing the development of more explainable, adaptable, and reliable AI systems, and contributing to the broader pursuit of Artificial General Intelligence (AGI).